8.Division by Zero is...
a.undefined
b.bounded
c.Infinite
d.finite

9.let \\(f(x)= 1/x)\\, then \\(f^{-1}(x))\\ is given as
a.x+1
b.none of the option
c.x
d.1/x
1
Expert's answer
2019-09-12T10:02:16-0400

9.let \\(f(x)= 1/x)\\, then \\(f^{-1}(x))\\ is given as


d.1/x


First you have to f (x) = y representing the output.


The function can be written as "y=\frac{1}{x}"


The variable "x" represents the input.


To find the inverse, the input is exchanged for the output, that is,


"x=\frac{1}{y}"


It is solved for "x":


"x=f^{(-1)}(y)=\frac{1}{y}\\"


Finally, we replace the variable.


The inverse function is "f^{-1}(X)=\frac{1}{X}"